OpenStudy (anonymous):

A copy of was made of sallys photograph the copy resized the photograph so that it measured 4 inches wide by 8 inches long if the actual photograph is 20 inches wide what is the length? what is the length? Choices 12 inches 32 inches 40 inches 48 inches @This_Is_Batman @texaschic101 @bibby

OpenStudy (anonymous):

See

OpenStudy (texaschic101):

is this the entire question....because it seems like something is missing

OpenStudy (anonymous):

the rest is...

OpenStudy (anonymous):

Whoa.. Something's not quite right here..

OpenStudy (anonymous):

It measured 4 inches wide by 8 inches long if the actual photograph is 20 inches wide what is the length?

OpenStudy (anonymous):

there

OpenStudy (anonymous):

My mom is rushing me so can i plz just get the answer?

OpenStudy (anonymous):

@texaschic101 Would this be cross multiplying?

OpenStudy (texaschic101):

4/8 = 20/L cross multiply 4L = (20)(8) 4L = 160 L = 160/4 L = 40 length is 40

OpenStudy (anonymous):

thank uuuuuuuuuuuuuuuuu

OpenStudy (texaschic101):

anytime :).....hey...give Batman a medal

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Tex u r a good sport

OpenStudy (anonymous):

And cool

OpenStudy (texaschic101):

thx...I try

OpenStudy (anonymous):

well thx i gtg now U GUYS ROCK

OpenStudy (texaschic101):

c ya later :)

OpenStudy (anonymous):

Sorry. I knew how to do it... I left to my question. Oops..